### v1

@@ Abstract
Current multimodal models leveraging contrastive learning often face limitations in developing fine-grained conceptual understanding. This is due to random negative samples during pretraining, causing almost exclusively very dissimilar concepts to be compared in the loss function. Consequently, the models struggle with fine-grained semantic differences. To address this problem, we introduce a novel pretraining method incorporating synthetic hard negative text examples. The hard negatives permute terms corresponding to visual concepts, leading to a more fine-grained visual and textual concept alignment. Further, we introduce InpaintCOCO, a new challenging dataset for assessing the fine-grained alignment of colors, objects, and sizes in vision-language models. We created the dataset using generative inpainting from COCO images by changing the visual concepts so that the images no longer match their original captions. Our results show significant improvements in fine-grained concept understanding across a wide range of vision-language datasets, including our InpaintCOCO dataset.
Keywords:
Vision-Language Contrastive Learning Hard Negatives
$I_{1}$
$T^{pos}_{1}$
$T^{neg}_{2}$
$T^{neg}_{3}$
$T^{hn}_{1}$
a white cat sits under a black open umbrella.
a person standing on a loading platform next to a train.
a man leans near the door of a creatively painted tour bus.
a brown cat sits under a black open umbrella.
Figure 1: Classical contrastive learning approaches use $(I_{1},T^{pos}_{1})$ as positive pairs in combination with negative samples like $T^{neg}_{2}$ and $T^{neg}_{3}$ to learn a image-text alignment. A bag of words (e.g., nouns) is often sufficient to extract the correct text match to a given image, resulting in only broad concepts learned. We also use hard negatives like $T^{hn}_{1}$ so that fine-grained semantic concepts are learned for visual and textual alignment.
@@ 1 Introduction
Recent advancements in vision-language (VL) modeling have demonstrated the effectiveness of contrastive learning in various multimodal tasks [13, 5, 24]. However, this training method does not provide sufficient training signals for several important visual concepts [29]. We attributed it to the objective function that uses random and, therefore, too dissimilar negative samples, which prevents the model from learning fine-grained semantic representations of the concepts.
Therefore, we propose a novel approach to address the issue of poorly represented concepts in contrastive learning. We introduce a mechanism to incorporate hard negative samples into the contrastive learning loss. Specifically, we generate synthetic hard negative samples by permuting keywords in the captions of original image-text pairs, disrupting the alignment between the image content and its description.
@@
  • (i)
    We present a novel method for using hard negative samples in the contrastive learning objective, allowing the model to focus on refining its understanding of concepts.
  • (ii)
    By introducing hard negative samples in the language component, we compel the model to learn proper visual and language alignment. Our approach improves multimodal performance, although it operates exclusively on the language side of the model.
  • (iii)
    To evaluate the model from the visual perspective, we create a challenge set with over 1,260 adversarial examples by using generative image inpainting. This dataset serves as a comprehensive benchmark, allowing us to assess the model’s ability to validate its conceptual understanding. This is because the image was created in a standardized setting in which only a small part was changed.
In this work, we conduct extensive evaluations on four basic concepts – color, object, location, and size. These concepts were selected as examples to demonstrate the effectiveness and robustness of our proposed approach in capturing nuanced semantic relations, but it is important to note that the choice of concepts is flexible and can be tailored to specific applications. Furthermore, our methodology is easy to construct, requiring only minimal domain expertise and the simple usage of regular expressions. This study bridges the gap in contrastive learning, demonstrating that simple tweaks can significantly enhance multimodal understanding and model performance
Figure 2: Hard negative contrastive learning: Keyword permuter produces hard negative text samples which are then randomly injected for each image $u_{i}$, replacing a simple negative sample in InfoNCE loss.
@@ 2 Vision-Language Representation Learning
The objective of contrastive representation learning is to learn representations that are close to each other for similar samples and distant from each other for dissimilar samples. While many objectives originally addressed a single modality [1, 17, 20], the idea can also be extended to multimodal training as well.
A successful example of multimodal learning is CLIP [13]. CLIP is a Transformer-based model that consists of an image encoder and a text encoder, which are trained simultaneously. The objective is to maximize the cosine similarity of the image and text embeddings from the correct image-text pairs and to minimize the similarity between the incorrect pairs. A batch of $N$ training samples (i.e., matching image-text pairs) results in a similarity matrix for each image-text combination. The main diagonal indicates the correct pair matches; the remaining entries correspond to negative entries. The InfoNCE loss [10, 27] is applied on $N\times N$ similarity scores.
This heuristical construction of negative samples has several issues. In some cases, negative captions can still match the given image, especially if the text is short and lacks details. Additionally, the negative pairs are often very dissimilar, which causes the model to decide only on coarse-grained features.
We illustrate this in Figure 1 with text-image pair $I_{1}$ and $T_{1}^{pos}$. The original learning process only uses negative text samples, such as $T^{neg}_{2}$ and $T^{neg}_{3}$. Here, it is sufficient that the model can assign or negate objects from texts (i.e., nouns) to the objects in the image. The fact that no “person” and no “door” are present in image $I_{1}$ is sufficient for the model to be able to discard these image descriptions. As a result, fine-grained concepts like object-color alignment, size, or spatial details (“cat under umbrella”) are not necessary for reaching low loss. The model can rely solely on the presence and absence of specific objects in this example. In this scenario, the caption can be seen as a bag of words without linguistic structure. We address this problem by creating new hard negative data samples to learn more fine-grained concepts. See § 3 for details.
@@
Radenovic et al. [12] use importance sampling (based on [14]), which up-samples hard negative samples and down-samples or ignores simple negatives. Yet, they reweight the simple negative samples per batch only but do not create new, challenging samples requiring fine-grained understanding.
Rösch and Libovický [16] propose keyword permutation to create hard negative samples to learn spatial concepts. They add a spatial understanding classifier as an auxiliary pretraining objective and evaluated the models on visual question answering (VQA). Their model is based on LXMERT [21] and not on a contrastive learning approach like CLIP.
Doveh et al. [2] generate hard negatives using a rule-based procedure where they permute keywords. They randomly parse parts of speech and fill the mask using a BERT encoder with a plausible but wrong word. Unlike us, they do not incorporate hard negatives into the similarity matrix but use an auxiliary loss that is added to the original contrastive loss. They use, in total, four distinct loss functions, introducing an additional layer of complexity to the overall procedure. In contrast, our approach adheres to the usage of the original loss function, with minimal modifications limited to the text inputs.
@@ 3 Contrastive Learning with Hard Negatives
We present a novel contrastive learning approach using sampled negative pairs and artificially generated textual hard negatives. Instead of training with one positive sample and several weak negative samples, we create a scenario where models also minimize the similarity to hard negative textual samples. This forces the model to learn fine-grained concepts during the training process.
We use keyword permutations for different concepts to break the correct meaning of an image caption. See Figure 1 for an example of the concept color. As a result, the new caption still lists, e.g., correct objects (e.g., “cat” and “umbrella”) and actions (e.g., “cat sits”) from the image but is no longer correct. We call this a “hard negative sample”. Using these samples during training, we ensure that fine-grained concepts are learned. The idea to inject hard negative samples in contrastive learning is highlighted in Figure 2. The InfoNCE loss (following [7]) is defined as
$\begin{split}\mathcal{L}_{\text{InfoNCE}}(\mathbf{U},\mathbf{V}):=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\mathbf{u}_{i}^{\top}\mathbf{v}_{i}/\tau)}{\sum_{j=1}^{N}\exp(\mathbf{u}_{i}^{\top}\mathbf{v}_{j}/\tau)}\\ -\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\mathbf{u}_{i}^{\top}\mathbf{v}_{i}/\tau)}{\sum_{j=1}^{N}\exp(\mathbf{u}_{j}^{\top}\mathbf{v}_{i}/\tau)},\end{split}$ (1)
where $\mathbf{u_{i}}$ is the learned visual embedding and $\mathbf{v_{i}}$ the textual embedding of image $I_{i}$ and caption $T_{i}$ respectively, and one or more $\mathbf{v_{j}}$ represent the embedding of the hard negative sample of $T_{i}$, which is denoted as $T_{i}^{hn}$ in Figure 1. $\mathbf{U}$ and $\mathbf{V}$ notes the representations within a batch of size $N$. The temperature is depicted as $\tau$.
@@ 3.1 Creating Hard Negative Text Samples
For various concepts, we replace specific keywords using a regex-based tool attached in the software submission. For example, we replace “white cat” with “brown cat” or “cat” with “dog”. We create permutation heuristics for four different concepts, namely colors, objects, size, location:
@@
  • For location keywords we use 12 one-to-one permutation relations.
  • For size keywords we use 11 one-to-one permutation relations.
The full list of heuristics is shown in the Supplementary. The created samples are used for training and for evaluation in § 5.1 and § 5.3.
@@ 3.2 Training Details
In our experiments, we use CLIP’s image and text encoder from the ViT-B/32 model, [^1] https://huggingface.co/openai/clip-vit-base-patch32 which has 87 million and 63 million parameters respectively. (However, our code is modular, and any Hugging Face image and text encoder can be used in the framework.) The code is based on transformers v4.30 [23].
We train all models using a batch size of 64. For concepts where several negative examples exist (i.e., color, object), we train models with 1, 2, and 3 hard negatives. This results in a proportion of hard negatives of 1.5%, 3%, and 4.5%, respectively. We only work with one hard negative for the other concepts.
@@ 3.3 Training Data
CLIP is pretrained on 400 million image-text pairs from web sources, and we continue the model pretraining. We use the COCO image captioning dataset [8] which has 591 thousand image-text pairs (2017 train version). For each concept, we filtered out samples where at least one keyword was present so that a keyword permutation could be applied. For evaluation, we use the validation set of COCO 2017. The respective dataset sizes are shown in the Supplementary.
@@ 4 InpaintCOCO: Challenge Set from the Visual Perspective
[>=latex, node distance=2cm]
(box1) [draw, rectangle, minimum width=2cm, minimum height=4.5cm, align=center] at (0,0) COCO sample
A red fire
hydrant is on
the sidewalk
next to a sign. ;
[->] (box1.east) – node[below,align=center,rotate=90,anchor=west] Mask
“fire hydrant” ++(1.05,0);
(box2) [draw, rectangle, minimum width=2cm, minimum height=3cm, right of=box1, node distance=3.3cm, align=center] Mask
Refer to caption ;
[->] (box2.east) – node[below,align=center,rotate=90,anchor=west]Inpaint “yellow
fire hydrant” ++(1.075,0);
(box3) [draw, rectangle, minimum width=2cm, minimum height=3cm, right of=box2, node distance=3.3cm, align=center] Inpainted
Refer to caption ;
[->] (box3.east) – node[below,align=center,rotate=90,anchor=west] Add Caption ++(.85,0);
(box4) [draw, rectangle, minimum width=2cm, minimum height=4.5cm, right of=box3, node distance=3.1cm, align=center] New sample
A yellow fire
hydrant is on
the sidewalk
next to a sign. ;
Figure 3: Creation of hard negative image samples using open vocabulary segmentation for the masking prompt and text-to-image generation for the inpainting prompt. Additionally, a new correct caption is created manually. The InpaintCOCO dataset was created for concepts object, color, and size.
Many multimodal tasks, such as VL Retrieval and Visual Question Answering, present results in terms of overall performance. Unfortunately, this approach overlooks more nuanced concepts, leaving us unaware of which specific concepts contribute to the success of current models and which are ignored. In response to this limitation, more recent benchmarks attempt to assess particular aspects of vision-language models. Some existing datasets focus on linguistic concepts utilizing one image paired with multiple captions; others adopt a visual or cross-modal perspective. In this study, we are particularly interested in fine-grained visual concept understanding, which we believe is not covered in existing benchmarks in sufficient isolation. Therefore, we create the InpaintCOCO dataset which consists of image pairs with minimum differences that lead to changes in the captions.
@@
Benchmarks such as ARO [26] or VL-CheckList [28] evaluate models from the language perspective. ARO examines understanding of attributes and relations without using specific concepts such as color or size. VL-CheckList is a dataset that investigates concrete concepts such as location, size, material, color, and relations, yet parts of the dataset are not available anymore.
On the other side, SVO [4] is a dataset that allows analysis from the visual perspective (2 images with 1 caption). Here, relations that deal with verbs are examined. To our knowledge, the only dataset that deals with fine-grained comprehension from a cross-model perspective is Winoground [22]. The dataset consists of two image-text pairs that are very similar to each other. This benchmark probes object relations that do not refer to specific concepts. The images shows similar concepts but are very dissimilar in terms of overall appearance. For samples of the two latter datasets, see the Supplementary. Both datasets contain real-world images that are in some ways similar in terms of objects, but the scenes still differ significantly. Therefore, it is difficult to tell which image differences cause the model predictions. We overcome this limitation by creating InpaintCOCO, the first dataset with only minor changes in the visual components, so that concept comprehension can be analyzed in a more standardized setting. An overview is listed in Table 1.
Dataset Size Perspective Samples Tasks
ARO 77k Linguistic 1 image $\leftrightarrow$ 2 texts Attributes, Relations, Order
VL-CheckList 410k Linguistic 1 image $\leftrightarrow$ 2 texts Attributes, Relations, Object
SVO-Probs 48k Visual 2 images $\leftrightarrow$ 1 text Relations (Verb underst.)
InpaintCOCO 1,260 Cross-modal 2 images $\leftrightarrow$ 2 texts Attributes, Object underst.
Table 1: Datasets for fine-grained understanding in VL.
@@
The dataset creation process can be viewed as complementary to textual hard-negative samples (§ 3) in the visual domain. Unlike keyword permutation, this cannot be done automatically with sufficient accuracy. Even though image segmentation and generative inpainting tools reach impressive results, they still require human supervision to produce high-quality images. Creating a high-quality test set therefore requires annotation work.
To generate hard negative image samples, we need to change individual details in the image so that the textual image description no longer fits. The procedure is illustrated in Figure 3. The annotation proceeds as follows: The annotators are provided with an image and decide if the image is suitable for editing. If yes, they input the prompt for the object that should be replaced. Using the open vocabulary segmentation model CLIPSeg [^2] https://huggingface.co/CIDAS/clipseg-rd64-refined [9] we obtain a mask for our object of interest (i.e., “fire hydrant”). Then, the annotator inputs a prompt for Stable Diffusion v2 Inpainting [^3] https://huggingface.co/stabilityai/stable-diffusion-2-inpainting [15] (e.g. with the prompt “yellow fire hydrant”), which shows three candidate images. The annotators can try new prompts or skip the current image if the result is insufficient. Finally, the annotator enters a new caption that matches the edited image. See the Supplementary for more detail. The images and captions come from the COCO 2017 validation data. We only use images that contain the desired concept and where licenses allow adaptations.
Object Color Size
Original
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
A bird is standing on top of a car.
A couple of white bathroom sinks sitting next to a toilet.
A living room with white furniture and a small wooden table.
Inpainted
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
A cat is standing on top of a car.
A couple of black bathroom sinks sitting next to a toilet.
A living room with white furniture and a huge wooden table.
Table 2: InpaintCOCO sample pairs for all concepts.
We provide 452 images for the concept object, 465 for color, and 343 for size. Some examples are displayed in Table 2. In contrast to the training process, objects in images are only replaced with objects from the same COCO super category, i.e., “cat” with another animal or “chair” for another piece of furniture. Since location would require erasing at one spot and implanting objects at another in a nontrivial way (especially regarding depth), we discard this one concept in the newly created dataset. The dataset is available via Hugging Face. [^4] https://huggingface.co/datasets/phiyodr/InpaintCOCO

### v2

@@ Abstract
Current vision-language models leveraging contrastive learning often face limitations in developing fine-grained conceptual understanding. This is due to random negative samples during pretraining, causing almost exclusively very dissimilar concepts to be compared in the loss function. Consequently, the models struggle with fine-grained semantic differences. To address this problem, we introduce a novel pretraining method incorporating synthetic hard negative text examples. The hard negatives replace terms corresponding to visual concepts, leading to a more fine-grained visual and textual concept alignment. Further, we introduce InpaintCOCO, a new challenging dataset for assessing the fine-grained alignment of colors, objects, and sizes in vision-language models. We created the dataset using generative inpainting from COCO images by changing the visual concepts so that the images no longer match their original captions. Our results show significant improvements in fine-grained concept understanding across various vision-language datasets, including our InpaintCOCO dataset.
Enhancing Conceptual Understanding in Multimodal Contrastive Learning through Hard Negative Samples
Philipp J. Rösch1, Norbert Oswald1, Michaela Geierhos1,  and Jindřich Libovický2 1Bundeswehr University Munich, Germany 2Faculty of Mathematics and Physics, Charles University, Czech Republic {philipp.roesch,norbert.oswald,michaela.geierhos}@unibw.de libovicky@ufal.mff.cuni.cz
$I_{1}$ Refer to caption $T^{pos}_{1}$ $T^{neg}_{2}$ $T^{neg}_{3}$ $T^{hn}_{1}$
a white cat sits under a black open umbrella. a person standing on a loading platform next to a train. a man leans near the door of a creatively painted tour bus. a brown cat sits under a black open umbrella.
Figure 1: Classical contrastive learning approaches use $(I_{1},T^{pos}_{1})$ as positive pairs in combination with negative samples like $T^{neg}_{2}$ and $T^{neg}_{3}$ to learn an image-text alignment. A bag of words (e.g., nouns) is often sufficient to extract the correct text that matches a given image, resulting in only broad concepts learned. We also use hard negatives like $T^{hn}_{1}$ so that fine-grained semantic concepts are learned for visual and textual alignment.
@@ 1 Introduction
Recent advancements in vision-language (VL) modeling have demonstrated the effectiveness of contrastive learning in various multimodal tasks (Radford et al., 2021; Jia et al., 2021; Yao et al., 2021). However, this training method does not provide sufficient training signals for several important visual concepts Zhao et al. (2023). We attributed it to the objective function’s use of random and, therefore, too dissimilar negative samples, which prevents the model from learning fine-grained semantic representations of the concepts.
Therefore, we propose a novel approach to address the issue of poorly represented concepts in contrastive learning. We introduce a mechanism to incorporate hard negative samples into the contrastive learning loss. Specifically, we generate synthetic hard negative samples by substituting keywords in the captions of original image-text pairs, disrupting the alignment between the image content and its description.
@@
(i) We present a novel method for using hard negative samples in the contrastive learning objective, allowing the model to focus on refining its understanding of concepts.
(ii) By introducing hard negative samples in the language component, we compel the model to learn proper visual and language alignment. Our approach improves multimodal performance, although it operates exclusively on the language side of the model.
(iii) To evaluate the model from the visual perspective, we create a challenge set with over 1,260 adversarial examples by using generative image inpainting. This dataset serves as a comprehensive benchmark, allowing us to assess the model’s ability to validate its conceptual understanding. This is because the image was created in a standardized setting in which only a small part was changed.
In this work, we conduct extensive evaluations of four basic concepts – color, object, location, and size. These concepts were selected as examples to demonstrate the effectiveness and robustness of our proposed approach in capturing nuanced semantic relations, but it is important to note that the choice of concepts is flexible and can be tailored to specific applications. Furthermore, our methodology is easy to construct, requiring only minimal domain expertise and the simple usage of regular expressions. This study shows that simple tweaks in contrastive learning can significantly enhance multimodal understanding and model performance.
Figure 2: Hard negative contrastive learning: Keyword substitution produces hard negative text samples, which are then randomly injected for each image $u_{i}$, replacing a simple negative sample in InfoNCE loss.
@@ 2 Vision-Language Representation Learning
The objective of contrastive representation learning is to learn representations that are close to each other for similar samples and distant from each other for dissimilar samples. While many objectives originally addressed a single modality (Chopra et al., 2005; Schroff et al., 2015; Sohn, 2016; Oord et al., 2018), the idea can also be extended to multimodal training as well.
A successful example of multimodal learning is CLIP (Radford et al., 2021). CLIP is a Transformer-based model that consists of an image encoder and a text encoder, which are trained simultaneously. The objective is to maximize the cosine similarity of the image and text embeddings from the correct image-text pairs and to minimize the similarity between the incorrect pairs. A batch of $N$ training samples (i.e., matching image-text pairs) results in a similarity matrix for each image-text combination. The main diagonal indicates the correct pair matches; the remaining entries correspond to negative entries. The symmetric cross-entropy loss is applied on $N\times N$ similarity scores.
This heuristical construction of negative samples has several issues. Negative captions can still match the given image in some cases, especially if the text is short and lacks details. Additionally, the negative pairs are often very dissimilar, which causes the model to decide only on coarse-grained features.
We illustrate this in Figure 1 with text-image pair $I_{1}$ and $T_{1}^{pos}$. The original learning process only uses negative text samples, such as $T^{neg}_{2}$ and $T^{neg}_{3}$. Here, it is sufficient that the model can assign or negate objects from texts (i.e., nouns) to the objects in the image. The fact that no “person” and no “door” are present in image $I_{1}$ is sufficient for the model to discard these image descriptions. As a result, fine-grained concepts like object-color alignment, size, or spatial details (“cat under umbrella”) are not necessary for reaching low loss. In this example, the model can rely solely on the presence and absence of specific objects. In this scenario, the caption can be seen as a bag of words without linguistic structure.
We address this problem by creating new hard negative data samples to learn more fine-grained concepts. See § 3 for details.
@@
Radenovic et al. (2023) use importance sampling (based on Robinson et al., 2021), which up-samples hard negative samples and down-samples or ignores simple negatives. Yet, they reweight the simple negative samples per batch only but do not create new, challenging samples requiring fine-grained understanding.
Rösch and Libovický (2022) propose keyword permutation to create hard negative samples to learn spatial concepts. They added a spatial understanding classifier as an auxiliary pretraining objective and evaluated the models on visual question answering (VQA). Their model is based on LXMERT (Tan and Bansal, 2019) and not on a contrastive learning approach like CLIP.
Doveh et al. (2023) generate hard negatives using a rule-based procedure where they replace keywords. Moreover, they implement an approach where they randomly parse parts of speech and fill the mask using a BERT encoder with a plausible but wrong word. Unlike us, they do not incorporate hard negatives into the similarity matrix but use an auxiliary loss summed with the original contrastive loss. They use four distinct loss functions in total, introducing an additional layer of complexity to the overall procedure. In contrast, our approach uses the original loss function, with minimal modifications limited to the text inputs.
@@ 3 Contrastive Learning with Hard Negatives
We present a novel contrastive learning approach using sampled negative pairs and artificially generated textual hard negatives.
Instead of training with one positive sample and several weak negative samples, we create a scenario where models also minimize the similarity to hard negative textual samples. This forces the model to learn fine-grained concepts during training.
We use keyword substitutions for different concepts to break the correct meaning of an image caption. See Figure 1 for an example of the concept color. As a result, the new caption still lists, e.g., correct objects (e.g., “cat” and “umbrella”) and actions (e.g., “cat sits”) from the image but is no longer correct. We call this a “hard negative sample”. Using these samples during training, we ensure that fine-grained concepts are learned. The idea to inject hard negative samples in contrastive learning is highlighted in Figure 2.
@@ 3.1 Creating Hard Negative Text Samples
COCO sampleRefer to captionA red fire hydrant is on the sidewalk next to a sign. Mask“fire hydrant”MaskRefer to captionInpaint “yellow fire hydrant”InpaintedRefer to captionAdd CaptionNew sampleRefer to captionA yellow fire hydrant is on the sidewalk next to a sign.
Figure 3: Create hard negative image samples using open vocabulary segmentation for the masking prompt and text-to-image generation for the inpainting prompt. Additionally, a new correct caption is created manually. The InpaintCOCO dataset was created for concepts object, color, and size.
For various concepts, we replace specific keywords using a regex-based tool. For example, we replace “white cat” with “brown cat” or “cat” with “dog”. We create substitution heuristics for four different concepts, namely colors, objects, size, location:
@@
  • For location keywords we use 12 one-to-one substitution relations.
  • For size keywords we use 11 one-to-one substitution relations.
The full list of heuristics is shown in Table 3 in the Appendix. The created samples are used for training and evaluation in § 5.1 and § 5.3.
The keywords were selected based on dataset statistics. For specific applications, a domain expert may have to select other terms (e.g. specific colors in the fashion industry).
@@ 3.2 Training Details
In our experiments, we use CLIP’s image and text encoder from the ViT-B/32 model, [^1] https://huggingface.co/openai/clip-vit-base-patch32 which has 87 million and 63 million parameters respectively. And hence, is a relatively small model compared with current multimodal GenAI models. Our code is modular, and any image and text encoder from transformers (Wolf et al., 2020) can be used in the framework.
We train all models using a batch size of 64. For concepts with multiple negative examples (i.e., color, object), we train models with 1, 2, and 3 hard negatives. This results in a proportion of hard negatives of 1.5%, 3%, and 4.5%, respectively. We only use one hard negative for the other concepts.
@@ 3.3 Training Data
CLIP is pretrained on 400 million image-text pairs from web sources, and we continue the model pretraining. We use the COCO image captioning dataset (Lin et al., 2015), which has 591 thousand image-text pairs (2017 train version). For each concept, we filtered out samples where at least one keyword was present so that a keyword substitution could be applied. For evaluation, we use the validation set of COCO 2017. The respective dataset sizes are shown in the Appendix in Table 5.
@@ 4 InpaintCOCO: Challenge Set from the Visual Perspective
Many multimodal tasks, such as VL Retrieval and Visual Question Answering, present results in terms of overall performance. Unfortunately, this approach overlooks more nuanced concepts, leaving us unaware of which specific concepts contribute to the success of current models and which are ignored. More recent benchmarks attempt to assess particular aspects of vision-language models in response to this limitation. Some existing datasets focus on linguistic concepts utilizing one image paired with multiple captions; others adopt a visual or cross-modal perspective. In this study, we are particularly interested in fine-grained visual concept understanding, which we believe is not covered in existing benchmarks in sufficient isolation. Therefore, we create the InpaintCOCO dataset with image pairs with minimum differences that lead to changes in the captions.
@@
Benchmarks such as ARO (Yuksekgonul et al., 2023) or VL-CheckList (Zhao et al., 2022) evaluate models from the language perspective. ARO examines understanding of attributes and relations without using specific concepts such as color or size. VL-CheckList [^2] Parts of the dataset are not available anymore. is a dataset that investigates concrete concepts such as location, size, material, color, and relations.
On the other side, SVO (Hendricks and Nematzadeh, 2021) is a dataset that allows analysis from the visual perspective (2 images with 1 caption). Here, relations that deal with verbs are examined. To our knowledge, the only dataset that deals with fine-grained comprehension from a cross-model perspective is Winoground (Thrush et al., 2022). The dataset consists of two image-text pairs that are very similar to each other. This benchmark probes object relations that do not refer to specific concepts. The images show similar concepts but are very dissimilar in overall appearance. For samples of the two latter datasets, see Figure 5 in the Appendix. Both datasets contain real-world images that are in some ways similar in terms of objects, but the scenes still differ significantly. Therefore, it is difficult to tell which image differences cause the model predictions.
We overcome this limitation by creating InpaintCOCO, the first dataset with only minor changes in the visual components, so that concept comprehension can be analyzed in a more standardized setting.
@@
The dataset creation process can be viewed a complement to textual hard-negative samples (§ 3) in the visual domain. Unlike keyword substitutions, this cannot be done automatically with sufficient accuracy. Even though image segmentation and generative inpainting tools reach impressive results, they still require human supervision to produce high-quality images. Creating a high-quality test set, therefore, requires annotation work.
To generate hard negative image samples, we need to change individual details in the image so that the textual image description no longer fits. The procedure is illustrated in Figure 3.
The annotation proceeds as follows: The annotators are provided with an image and decide if they want to edit it. If yes, they input the prompt for the object that should be replaced. Using the open vocabulary segmentation model CLIPSeg [^3] https://huggingface.co/CIDAS/clipseg-rd64-refined Lüddecke and Ecker (2022) we obtain a mask for our object of interest (i.e., “fire hydrant”). Then, the annotator inputs a prompt for Stable Diffusion v2 Inpainting [^4] https://huggingface.co/stabilityai/stable-diffusion-2-inpainting Rombach et al. (2022) (e.g. with the prompt “yellow fire hydrant”), which shows three candidate images. The annotators can try new prompts or skip the current image if the result is insufficient. Finally, the annotator enters a new caption that matches the edited image. See Appendix A for all details. The images and captions come from the COCO 2017 validation data. We only use images that contain the desired concept and where licenses allow adaptations.
We provide 452 images for the concept object, 465 for color, and 343 for size. In contrast to the training process, objects in images are only replaced with objects from the same COCO super category, i.e., “cat” with another animal or “chair” for another piece of furniture. Since location would require erasing at one spot and implanting objects at another in a nontrivial way (especially regarding depth), we discard this one concept in the newly created dataset. The dataset will be available upon publishing via the HuggingFace hub.